A certain cross between pea plants produces progeny in the ratio: 4 yellow and 2 green, if 10 progenies are chosen random, find proability;

a. Those 10 chosen are yellow?

b Exactly 6 out of 10 chosen are green?

c. What is the probability that 5 out of the 10 chosen are of the same color? Could be (5 green) or (5 yellow).
Answer #1

a) here for binomial distribuion probability of yellow =4/6 =2/3 and n=10

P(Those 10 chosen are yellow) =(2/3)10 =0.0173

b)

for binomial distribuion probability of green =2/6 =1/3 and n=10

P(Exactly 6 out of 10 chosen are green)=\binom{10}{6}(1/3)^{6}(2/3)^{4} =0.0569

c)

probability that 5 out of the 10 chosen are of the same color=P(5 of yellow color and 5 green)

=\binom{10}{5}(1/3)^{5}(2/3)^{5} =0.1366